# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 3.3 Application to $d$-regular Ramanujan graphs
$Q^{n}=\frac{1}{N}1\otimes 1+Q^{n}|_{\boldsymbol{1}^{\perp}}$
where $\boldsymbol{1}\otimes\boldsymbol{1}$ is the $N\times N$ matrix which entries are all equal to $1$. Setting
$\rho:=\max\{|\mu|;\leavevmode\nobreak\ \mu\in\operatorname{Sp}(Q);\leavevmode\nobreak\ |\mu|\neq 1\}$
@@
Besides for all $d$, there exists $c_{d}>0$ such that we have the following Alon-Boppana lower bound
@@
For the previous bound one can refer to [Moh10]. From the previous bound we have that for all integer $d$ and $10<c<$ 1 we have $\rho\geq c\frac{2\sqrt{d-1}}{d}$ for $N$ large enough and therefore for all $n\leq\frac{\ln(N)}{7\ln(d)}$ we have $\rho^{n}\geq N^{-1/38}$. Owing to (14) and (1) we have
$\kappa(Q)\leq\left(1+o_{N}(1)\right)\rho.$
@@
Therefore for any undirected $d$-regular graph verifying that $\rho<\frac{1}{2}$, for $N$ large enough we have $\rho_{n}=\mathrm{N}_{n}$. In particular the spectral radius of either $A$ or $A_{r}$ is upper bounded by the spectral radius of $a$ and $p$ respectively. As an example, we can consider Ramanujan graphs which by definition verify
$\rho\leq\frac{2\sqrt{d-1}}{d}$
@@
and therefore for all $d\geq 15$ we have $\rho<\frac{1}{2}$. An explicit example of such graphs are the Ramanujan graphs $X^{p,q}$ of degree $p+1$ and cardinality given by $\frac{q(q^{2}-1)}{2}$ where $(p,q)$ are prime numbers. Their construction is given in [DSV03, Chapter 4].
@@ 4 Intermediate construction
We give here a constructive definition of the operator valued probability space $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi)$ and the operators $(q,u)\in\mathcal{A}$ which verifies the first statement of 2.2, i.e.
  • $q$ has same $\ast$-distribution as $Q$,
@@
In a second time, we study the $\ast$-distribution of $a=u+q$ and $p=(1-r)u+rq$ and give an approximation of the spectral radius of $a$ and $p$.
@@ 4.1 Construction of $q$ and $u$ as adjacency operators of graphs
The construction given here is inspired by the quasi-tree construction described by Hermon, Sly and Sousi [HSS20]. We will consider $q$ and $u$ as adjacency operators of explicit graphs, meaning that we will give explicit writings of these operators allowing us to then compute $a$ and $p$ moments. One has to consider the construction of these graphs as follows. First, we take infinite and countable copies of the vertex set $V=[N]$ on which the adjacency operator $Q$ is defined. Then we consider each copy as a vertex and we put a tree structure on the set of copies. The tree is regular and the degree depends on the dimension $N$ of the operator $Q$. More precisely, we consider one of the copies as the origin (denoted by $H_{\emptyset}$ in the Figure 3). We consider that the origin has $N$ descendants which are $N$ other copies of the vertex set $[N]$. To each vertex $x$ of the origin we assign uniformly at random one of the $N$ descendant copies. In that descendant, we take uniformly at random one of its vertex $x^{\prime}$ to receive the ascending edge. The edge $(x,x^{\prime})$ is represented in blue in Figure 3 and the adjacency operator that will correspond to the graph defined by the blue edges (after constructing them all) is $u$. We repeat this operation for each descendant, taking at random the vertex set copies among the copies never seen before, and so on. The following Figure 3 shows an example for $N=3$.
@@
We give simultaneously the construction $u^{\ast}$, inverse of the unit operator $u$. To each blue edge we connect an inverse red edge, which connects it to its antecedent. We also construct the symmetric graph above the origin, where now the red edges corresponding to $u^{\ast}$ are drawn first and can be seen as descendants, and the blue ones corresponding to $u$ are their inverse (see Figure 4).
@@
Finally it remains to place on each set of vertices the edges corresponding to the graph of $Q$ (in Figure 5 we considered $Q=(1,1,1)^{\top}\cdot(1,1,1)$).
@@
The operators $q$ and $u$ we are looking for are the adjacency operators of the graphs described above. For instance, the operator $a=u+q$ is the adjacency operator of the graph where we considered the blue and green edges of Figure 5.
@@ 4.2 Construction of the underlying Hilbert space
$\mathcal{S}:=\{(k_{0},\epsilon_{1},k_{1},\epsilon_{2},k_{2},...,\epsilon_{n-1},k_{n-1},\epsilon_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ ;n\geq 1\leavevmode\nobreak\ \leavevmode\nobreak\ ;\forall t\leq n,\leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon_{t}\in\{+,-\}\text{ and }k_{t}\in[N]\}.$
Let $\{H_{\emptyset}\}\cup\{H_{s}\}_{s\in\mathcal{S}}$ be independent copies of $\mathbb{C}^{N}$. With respect to the previous Section 4.1, $\mathcal{S}$ is the index set for the countable and infinite copies of $[N]$ considered as $Q$ vertex set. One can think of $n$ as the future number of generations from the origin and $\epsilon_{t}$ as either descending from generation $t$ to generation $t+1$ taking $u$ or $u^{\ast}$. For $s:=(k_{0},\epsilon_{1},k_{1},...,k_{n-1},\epsilon_{n})\in\mathcal{S}$, we set the Hilbert space
@@
$\{e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{(k_{0},i_{1},...,i_{n-1},\epsilon_{n})}\leavevmode\nobreak\ ;\leavevmode\nobreak\ s=(k_{0},i_{1},...,i_{n},\epsilon_{n})\in\mathcal{S}\}.$
@@
$\mathcal{V}_{1}:=\{0\}\cup\{(k_{0},i_{1},i_{2},...,i_{n})\leavevmode\nobreak\ ;\leavevmode\nobreak\ n\in\mathbb{N}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \forall t\leq n,\leavevmode\nobreak\ i_{t}:=(\epsilon_{t},k_{t})\in\{+,-\}\times[N]\leavevmode\nobreak\ ;\text{ and }k_{0}\in[N]\}.$
@@
In particular, we see that $\mathrm{H}_{1}$ is equal to the closure of ${\operatorname{Span}\{e(v)\leavevmode\nobreak\ ;\leavevmode\nobreak\ v\in\mathcal{V}_{1}\}}$ and this space can be seen as all possible ways to connect any $n$-th generation vertex $v=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}$ to any descendant one $v^{\prime}:=(k_{0},i_{1},...,i_{n},i_{n+1})$ (note that $(v^{\prime})^{-}=v$). So for $e=e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{1}$ we can also denote:
@@
Recall that, with respect to the construction in Section 4.1, the randomness comes from the fact that at each generation, for each vertex $v$ we randomly take a descendant copy of the vertex set $[N]$ and a vertex $v^{\prime}$ in it to which we associate $v$. It will verify $(v^{\prime})^{-}=v$. Note that for $s:=(k_{0},i_{1},...,i_{n-1},\epsilon_{n})\in\mathcal{S}$ and $k\in[N]$, we have that $v:=(s,k)=(k_{0},i_{1},...,\epsilon_{n},k)\in\mathcal{V}_{1}$. So, keeping the same notations, from now on we denote
@@
$\mathcal{V}:=\{v=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\ell<n:\leavevmode\nobreak\ k_{\ell}\neq j_{-\epsilon_{\ell}}(k_{0},i_{1},...,i_{\ell-1})\text{ or }\epsilon_{\ell+1}=\epsilon_{\ell}\},$
@@
$\Xi:=\{e=e(k)\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ v_{e}\in\mathcal{V}\}.$ (15)
We denote by $\mathrm{H}$ the Hilbert space obtained. We now explicit the Banach and operator-valued probability space structure of $\mathcal{B}(\mathrm{H})\subset\mathcal{B}(L^{2}(\mathrm{H}_{0}))$.
@@
    $\forall h,h^{\prime}\in\mathrm{H},\leavevmode\nobreak\ \leavevmode\nobreak\ (h,h^{\prime}):=\mathbb{E}[\langle h,h^{\prime}\rangle]\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \|{h}\|_{2}:=\sqrt{(h,h)}.$
[…]
@@
    $(\mathcal{B}(\mathrm{H}),\Phi,\ast)$ is non-commutative probability space where the trace is given by
@@
    for $a\in\mathcal{B}(\mathrm{H})$, where $e_{\alpha}\in H_{\emptyset}$ and $\alpha\sim\mathcal{U}([N])$ is independent from $(j_{\epsilon}(s))_{\epsilon\in\{-,+\},s\in\mathcal{S}}$.
@@
    The conditional expectation $\mathrm{E}$ on $\mathcal{B}(\mathrm{H})$ is the operator verifying
@@
    for all $a\in\mathcal{B}(\mathrm{H})$ and for all $e,f\in\Xi$.
We now define the operators $(q,u)$ in 2.2.
@@
$\Pi$ $:\mathrm{M}_{N}(\mathbb{C})\longrightarrow\mathcal{B}(\mathrm{H})$
[…]
and we set $q:=\Pi(Q)$. We now consider the operator valued probability space $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi,\ast)$ where $\mathcal{A}$ is the subalgebra of $\mathcal{B}(\mathrm{H})$ spanned by $u$, $u^{\ast}$, $\mathcal{M}:=\Pi(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ and $\mathcal{B}:=\mathrm{E}(\mathcal{A})$. Finally we denote $\mathcal{U}$ the sub algebra of $\mathcal{A}$ spanned by $u$ and $u^{\ast}$.
@@
We could have defined both $u$ and $u^{*}:=u^{-1}$ as follows:

### v4

@@ 4.3 Application to $d$-regular Ramanujan graphs
$Q^{n}=\frac{1}{N}\mathbf{1}^{\top}\mathbf{1}+Q^{n}|_{\boldsymbol{1}^{\perp}}$
where $\mathbf{1}^{\top}\mathbf{1}$ is the $N\times N$ matrix which entries are all equal to $1$. Setting
$\rho:=\max\{|\mu|;\penalty 10000\ \mu\in\operatorname{Sp}(Q);\penalty 10000\ |\mu|\neq 1\}$
@@
Besides for all $d$, there exists $c_{d}>0$ such that we have the following Alon-Boppana lower bound [MOH10]
@@
From the previous bound we have that $\rho\geq c\frac{2\sqrt{d-1}}{d}$ for all integer $d$, $0<c<1$ and $N$ large enough. Therefore for all $n\leq\frac{\ln(N)}{7\ln(d)}$ we have $\rho^{n}\geq N^{-1/38}$. Owing to (13) and (1) we have
$\kappa(Q)\leq\left(1+o_{N}(1)\right)\rho.$ (14)
@@
In particular, for any undirected $d$-regular graph verifying that $\rho<\frac{1}{2}$, for $N$ large enough we have $\rho_{n;\mathrm{x}}=\mathrm{N}_{n,\mathrm{x}}$ ans the spectral radius of $A_{\mathrm{x}}$ is upper bounded by the spectral radius of $\mathrm{b}_{\mathrm{x}}$. As an example, we can consider Ramanujan graphs which by definition verifies
$\rho\leq\frac{2\sqrt{d-1}}{d}.$
@@
In that case, for all $d\geq 15$ we have $\|Q_{\perp}\|=\rho<\frac{1}{2}$, for $N$ large enough and owing to (14) we have $\kappa(Q)<1$ and therefore the upper bound for $|\lambda_{2}(A_{\mathrm{s}})|$ given by ˜2.1 is ${\mathrm{N}}_{n,\mathrm{s}}\geq 1$. An explicit example of such graphs are the Ramanujan graphs $X^{p,q}$ of degree $p+1$ and cardinality given by $\frac{q(q^{2}-1)}{2}$ where $(p,q)$ are prime numbers. Their construction is given in [DSV03, Chapter 4].
5 Overview of the proofs
The proof of ˜2.1 is independent of the rest of this paper. We give an explicit constructive definition of the free with amalgamation operators $(q,u)\in(\mathcal{A},\mathcal{B},\mathrm{E},,\Phi,\ast)$ in Section˜6. This construction is inspired by the existing construction given by Hermon, Sly and Sousi [HSS20] (see Figure˜3, Figure˜4 and Figure˜5). This construction allows us to compute the moments of the limit operators given by (9) (see Section˜6.4) and therefore the approximation of radius of the limit operators.
In Section˜7 we start by using that for any matrix $A\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ and integer $\ell\geq 1$ we have
$\rho(A)\leq\|A^{\ell}\|^{1/\ell}$
to obtain the bound (19) and we then state the main technical statement of this paper given by ˜7.1. In this statement we consider any power $\ell\geq 1$ and denoting $\mathbb{P}:=\mathbb{P}\left(|\lambda_{2}(A_{\mathrm{x}})|\geq(1+\epsilon)\rho_{n,\mathrm{x}}\right)$ we rewrite the probability we want to bound as follows
[…]
The notion of a $(M,Q)$ being $\ell$-tangle-free is introduced by ˜7.2. The couple $(M,Q)$ being $\ell$-tangle free tells us that looking at the powers of $A_{\mathrm{x}}$, the second probability in the sum above tends to zero with speed $N^{-c/2}$ for some $0<c<1$ provided that we have
$\ell<\frac{1-c}{3\ln(d)}\ln(N).$
The previous statement is the object of Section˜7.3 and ˜7.2. Once we showed that our couple $(M,Q)$ is $\ell$-tangle free with high probability, we rewrite the entries of $A^{\ell}_{\mathrm{x}}$ on this event (see (21)). Finally, we want to apply the property of random permutation given by [BOR15, Proposition 11], recalled in Section˜7.3. Owing to this proposition, we can define a new matrix $\underline{A}_{\mathrm{x}}^{(\ell)}$ where now all the appearance of $M$ entries are centered (see (26)). Indeed, introducing $\underline{A}_{\mathrm{x}}^{(\ell)}$, rest terms appear (see ˜7.1), and the normalized sum of these terms can be neglected.
In regard of ˜7.1 and assuming the operator norm of the rest terms are suitably controlled, it remains to upper bound the operator norm of $\underline{A}^{(\ell)}_{\mathrm{x}}$. This upper bound is given by Section˜8.1 and its proof, constituting the main technical challenge, is the object of Section˜8. In that section, the goal is to show that for an explicit power $\ell\sim\ln(N)$ we have
$\|\underline{A}_{\mathrm{x}}^{(\ell)}\|\leq\left(1+o(1)\right)^{\ell}\rho_{n,\mathrm{x}}^{\ell}$
with high probability and where $\rho_{n,\mathrm{x}}$ is introduced in (11). We will apply Markov inequality, i.e
$\mathbb{P}\left(\|\underline{A}_{\mathrm{x}}^{(\ell)}\|\leq\left(1+o(1)\right)^{\ell}\rho_{n,\mathrm{x}}^{\ell}\right)=\mathbb{P}\left(\|\underline{A}_{\mathrm{x}}^{(\ell)}{\underline{A}_{\mathrm{x}}^{(\ell)}}^{\top}\|^{m}\leq\left(1+o(1)\right)^{2\ell m}\rho_{n,\mathrm{x}}^{2m\ell}\right)\leq\frac{\mathbb{E}\left(\|\underline{A}_{\mathrm{x}}^{(\ell)}{\underline{A}_{\mathrm{x}}^{(\ell)}}^{\top}\|^{m}\right)}{\left(1+o(1)\right)^{2\ell m}\rho_{n,\mathrm{x}}^{2m\ell}}$
and bound the expectation in the left hand side of the inequality above using trace method (see the computation in Section˜8.1). The fact that we restrict to the event where $(M,Q)$ is $\ell$-tangle free allows us to rewrite the trace $\operatorname{Tr}\left\{\left(\underline{A}^{(\ell)}{\underline{A}^{(\ell)}}^{\top}\right)^{m}\right\}$ as a sum over all concatenations of $2m$ $\ell$-tangle free paths with specific bound properties (see (29)). We then define equivalent classes on the set of such $\ell$-tangle free paths and obtain the upper bound given by (33). We bound separately the number of equivalent classes and, for a given equivalent class, the contribution of the permutation matrix $M$ given by $\mu(\gamma)$ and the contribution of $Q$. Section˜8.2 and Section˜8.3 are dedicated to the proof of the bounds on respectively the cardinal of equivalent classes given by ˜8.1 and the contribution of the matrix $Q$ given by ˜8.4. The contribution of the permutation given by ˜8.7 is handled applying Section˜7.3. The strategy to encode and enumerate the number of equivalent class is analog to the various other high trace method proofs (see for instance [BOR15, BQZ18, BC19, BDH22]). We also conclude the proof of Section˜8.1 in Section˜8.3.
Finally, in Section˜8.4 we bound the rest terms appearing in ˜7.1 using the same strategy used to bound the operator norm of $\underline{A}_{\mathrm{x}}^{(\ell)}$. Indeed bounding the operator norms with traces we obtain (47), analog of (33) with a new definition for equivalent classes for paths. Section˜8.4 and ˜8.8 are the analog of Section˜8.1 and ˜8.1. The normalized sum in ˜7.1 makes the upper bound given in Section˜8.4 less challenging (see (50)). We gather all these results in Section˜9 to conclude the proof of ˜7.1.
@@ 6 Intermediate construction
We give here a constructive definition of the operator valued probability space $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi)$ and the operators $(q,u)\in\mathcal{A}$ which verifies the first statement of ˜2.1, i.e.
  • $q$ has, asymptotically, same $\ast$-distribution as $Q$,
@@
In a second time, we study the $\ast$-distribution of $\mathrm{b}_{\mathrm{s}}=u+q$ and $\mathrm{b}_{\mathrm{c}}=(1-r)u+rq$ and give an approximation of the spectral radius of $\mathrm{b}_{\mathrm{x}}$ for $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$.
@@ 6.1 Construction of $q$ and $u$ as adjacency operators of graphs
The construction given here is inspired by the quasi-tree construction described by Hermon, Sly and Sousi [HSS20]. We will consider $q$ and $u$ as adjacency operators of explicit graphs, meaning that we will give explicit writings of these operators allowing us to then compute $a$ and $p$ moments. One has to consider the construction of these graphs as follows. First, we take infinite and countable copies of the vertex set $V=[N]$ on which the adjacency operator $Q$ is defined. Then we consider each copy as a vertex and we put a tree structure on the set of copies. The tree is regular and the degree depends on the dimension $N$ of the operator $Q$. More precisely, we consider one of the copies as the origin (denoted by $H_{\emptyset}$ in the Figure˜3). We consider that the origin has $N$ descendants which are $N$ other copies of the vertex set $[N]$. To each vertex $x$ of the origin we assign uniformly at random one of the $N$ descendant copies. In that descendant, we take uniformly at random one of its vertex $x^{\prime}$ to receive the ascending edge. The edge $(x,x^{\prime})$ is represented in blue in Figure˜3 and the adjacency operator that will correspond to the graph defined by the blue edges (after constructing them all) is $u$. We repeat this operation for each descendant, taking at random the vertex set copies among the copies never seen before, and so on. The following Figure˜3 shows an example for $N=3$.
@@
We give simultaneously the construction $u^{\ast}$, inverse of the unit operator $u$. To each blue edge we connect an inverse red edge, which connects it to its antecedent. We also construct the symmetric graph above the origin, where now the red edges corresponding to $u^{\ast}$ are drawn first and can be seen as descendants, and the blue ones corresponding to $u$ are their inverse (see Figure˜4).
@@
Finally it remains to place on each set of vertices the edges corresponding to the graph of $Q$ (in Figure˜5 we considered $Q=(1,1,1)^{\top}\cdot(1,1,1)$).
@@
The operators $q$ and $u$ we are looking for are the adjacency operators of the graphs described above. For instance, the operator $a=u+q$ is the adjacency operator of the graph where we considered the blue and green edges of Figure˜5.
@@ 6.2 Construction of the underlying Hilbert space
$\mathcal{S}:=\{(k_{0},\epsilon_{1},k_{1},\epsilon_{2},k_{2},...,\epsilon_{n-1},k_{n-1},\epsilon_{n})\penalty 10000\ \penalty 10000\ ;n\geq 1\penalty 10000\ \penalty 10000\ ;\forall t\leq n,\penalty 10000\ \penalty 10000\ \epsilon_{t}\in\{+,-\}\text{ and }k_{t}\in[N]\}.$
Let $\{H_{\emptyset}\}\cup\{H_{s}\}_{s\in\mathcal{S}}$ be independent copies of $\mathbb{C}^{N}$. With respect to the previous Section˜6.1, $\mathcal{S}$ is the index set for the countable and infinite copies of $[N]$ considered as $Q$ vertex set. One can think of $n$ as the future number of generations from the origin and $\epsilon_{t}$ as either descending from generation $t$ to generation $t+1$ taking $u$ or $u^{\ast}$. For $s:=(k_{0},\epsilon_{1},k_{1},...,k_{n-1},\epsilon_{n})\in\mathcal{S}$, we set the Hilbert space
@@
$\{e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{(k_{0},i_{1},...,i_{n-1},\epsilon_{n})}\penalty 10000\ ;\penalty 10000\ s=(k_{0},i_{1},...,i_{n},\epsilon_{n})\in\mathcal{S}\}.$
@@
$\mathcal{V}_{1}:=\{0\}\cup\{(k_{0},i_{1},i_{2},...,i_{n})\penalty 10000\ ;\penalty 10000\ n\in\mathbb{N}\penalty 10000\ ;\penalty 10000\ \forall t\leq n,\penalty 10000\ i_{t}:=(\epsilon_{t},k_{t})\in\{+,-\}\times[N]\penalty 10000\ ;\text{ and }k_{0}\in[N]\}.$
@@
In particular, we see that $\mathrm{H}_{1}$ is equal to the closure of $\operatorname{Span}\{e(v)\penalty 10000\ ;\penalty 10000\ v\in\mathcal{V}_{1}\}$} and this space can be seen as all possible ways to connect any $n$-th generation vertex $v=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}$ to any descendant one $v^{\prime}:=(k_{0},i_{1},...,i_{n},i_{n+1})$ (note that $(v^{\prime})^{-}=v$). So for $e=e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\in\mathrm{H}_{1}$ we can also denote
@@
Recall that, with respect to the construction in Section˜6.1, the randomness comes from the fact that at each generation, for each vertex $v$ we randomly take a descendant copy of the vertex set $[N]$ and a vertex $v^{\prime}$ in it to which we associate $v$. It will verify $(v^{\prime})^{-}=v$. Note that for $s:=(k_{0},i_{1},...,i_{n-1},\epsilon_{n})\in\mathcal{S}$ and $k\in[N]$, we have that $v:=(s,k)=(k_{0},i_{1},...,\epsilon_{n},k)\in\mathcal{V}_{1}$. So, keeping the same notations, from now on we denote
@@
$\mathcal{V}:=\{v=(k_{0},i_{1},...,i_{n})\in\mathcal{V}_{1}\penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \forall\ell<n:\penalty 10000\ k_{\ell}\neq j_{-\epsilon_{\ell}}(k_{0},i_{1},...,i_{\ell-1})\text{ or }\epsilon_{\ell+1}=\epsilon_{\ell}\},$
@@
$\Xi:=\{e=e(k)\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})\penalty 10000\ \penalty 10000\ ;\penalty 10000\ v_{e}\in\mathcal{V}\}.$ (15)
We denote by $\mathrm{H}$ the Hilbert space obtained. We now explicit the Banach and operator-valued probability space structure of $\mathrm{B}(\mathrm{H})\subset\mathcal{B}(L^{2}(\mathrm{H}_{0}))$.
@@
    $\forall h,h^{\prime}\in\mathrm{H},\penalty 10000\ \penalty 10000\ (h,h^{\prime}):=\mathbb{E}[\langle h,h^{\prime}\rangle]\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \|{h}\|_{2}:=\sqrt{(h,h)}.$
    The subordinate norm $\|\cdot\|$ on $\mathcal{B}(\mathrm{H})$ gives to $\mathcal{B}(\mathrm{H})$ a Banach algebra structure.
@@
    $(\mathrm{B}(\mathrm{H}),\Phi,\ast)$ is non-commutative probability space where the trace is given by
@@
    for $a\in\mathrm{B}(\mathrm{H})$, where $e_{\alpha}\in H_{\emptyset}$ and $\alpha\sim\mathcal{U}([N])$ is independent from $(j_{\epsilon}(s))_{\epsilon\in\{-,+\},s\in\mathcal{S}}$.
@@
    The conditional expectation $\mathrm{E}$ on $\mathrm{B}(\mathrm{H})$ is the operator verifying
@@
    for all $a\in\mathrm{B}(\mathrm{H})$ and for all $e,f\in\Xi$.
We now define the operators $(q,u)$ in ˜2.1.
@@
$\Pi$ $:\mathrm{M}_{N}(\mathbb{C})\longrightarrow\mathrm{B}(\mathrm{H})$
$B$ $\longmapsto\left[\begin{aligned} b:\mathrm{H}&\rightarrow\mathrm{H}\\ e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes e(i_{n})&\mapsto e(k_{0})\otimes e(i_{1})\otimes\cdots\otimes Be(i_{n})\end{aligned}\right]$
and we set
${q}:=\Pi(Q_{\perp}).$ (16)
We now consider the operator valued probability space $(\mathcal{A},\mathcal{B},\mathrm{E},\Phi,\ast)$ where $\mathcal{A}$ is the subalgebra of $\mathrm{B}(\mathrm{H})$ spanned by $u$, $u^{\ast}$, $\mathcal{M}:=\Pi(\mathrm{M}_{N}(\operatorname{\mathbb{C}}))$ and $\mathcal{B}:=\mathrm{E}(\mathcal{A})$. Finally we denote $\mathcal{U}$ the sub algebra of $\mathcal{A}$ spanned by $u$ and $u^{\ast}$.
@@
We could have defined both $u$ and $u^{*}:=u^{-1}$ as follows